Question 1. Cost concepts The concepts of marginal cost (MC) and average cost (AC) are routinely used in economics. (a) S
Answer #1

ca) = 902892 (mcg) 0.59 AC (A) = 9+ 0.259 (6) Returns to scale when the cost of from de -increase in aty then increasing retuda TR - Pq Tha a +8.2592 - Pq - 9.8 6.2592 dt- & à 0.59 -0 (9= 20 AC = a + p = 6 + 0.289) Combination of increasing & decreas